### v1

# DexVLA: Vision-Language Model with Plug-In Diffusion Expert for General Robot Control

Abstract
Enabling robots to perform diverse tasks across varied environments is a central challenge in robot learning. While vision-language-action (VLA) models have shown promise for generalizable robot skills, realizing their full potential requires addressing limitations in action representation and efficient training. Current VLA models often focus on scaling the vision-language model (VLM) component, while the action space representation remains a critical bottleneck. This paper introduces DexVLA, a novel framework designed to enhance the efficiency and generalization capabilities of VLAs for complex, long-horizon tasks across diverse robot embodiments. DexVLA features a novel diffusion-based action expert, scaled to one billion parameters, designed for cross-embodiment learning. A novel embodiment curriculum learning strategy facilitates efficient training: (1) pre-training the diffusion expert that is separable from the VLA on cross-embodiment data, (2) aligning the VLA model to specific embodiments, and (3) post-training for rapid adaptation to new tasks. We conduct comprehensive experiments across multiple embodiments, including single-arm, bimanual, and dexterous hand, demonstrating DexVLA’s adaptability to challenging tasks without task-specific adaptation, its ability to learn dexterous skills on novel embodiments with limited data, and its capacity to complete complex, long-horizon tasks using only direct language prompting, such as laundry folding. In all settings, our method demonstrates superior performance compared to state-of-the-art models like Octo, OpenVLA, and Diffusion Policy.
I Introduction
Enabling robots to perform diverse tasks across varied environments is a central challenge in robotics. Achieving versatility—the ability to solve a variety of tasks across diverse environments, while adapting to language commands, environmental constraints, and unexpected disruptions—is even more demanding. Imitation learning [26, 40, 24, 59, 58, 35], particularly through vision-language-action (VLA) models (e.g., [5, 7, 23, 20]), has shown promise in enabling generalizable skills.
However, realizing the vision of omnipotent robot foundation models faces persistent challenges. Two key bottlenecks hinder progress: 1) Data scarcity: State-of-the-art models, like OpenVLA [23]/Octo [30], rely on massive datasets like the Open-X Embodiment dataset (4,000 hours) [31] or even larger corpora like the 10,000-hour dataset used by $\pi_{0}$ [5]. Collecting such data through human demonstrations is extremely costly and labor-intensive [13, 18, 51, 53]. (2) Architectural imbalance: current VLA models often prioritize scaling the vision-language model (VLM) component, i.e., OpenVLA uses 7B VLM and $\pi_{0}$ uses 3B VLM. Despite its enhanced visual and linguistic understanding through internet-scale data pretraining, the VLM component remains disconnected from the embodied, sensorimotor context of robotic action.
This paper introduces plug-in Diffusion Expert for Vision-Language-Action models, namely DexVLA, a novel framework designed to enhance the efficiency and generalization capabilities of VLA for complex, long-horizon tasks across diverse robot embodiments. We achieve this through two key innovations:
1) Billion-Parameter Diffusion Expert: Recognizing the limitations of conventional action experts, particularly in handling cross-embodiment data, we propose a new diffusion-based action expert. The diffusion expert utilizes a multi-head architecture, with each head corresponding to a specific embodiment, enabling effective learning across diverse morphologies. Furthermore, we scale the model size of the diffusion expert to one billion parameters, a substantial increase from the conventional multi-million parameter scale. This scaling significantly enhances the model’s capacity to learn intricate motor skills and control policies from diverse and extensive data.
2) Embodied Curriculum Learning: A three-stage training strategy that progressively learns harder tasks. This is conceptually similar to how human learn, which starts with simple tasks, and then gradually introduces complexity to avoid overwhelming the learner.
Stage 1: The cross-embodiment pre-training stage focuses on learning low-level, embodiment-agnostic motor skills. In this stage, we pre-train only the diffusion expert using cross-embodiment data, without involving the vision-language models.
Stage 2: The embodiment-specific alignment is a stage that analogy to ”adapt to your body”. Specifically, it bridges abstract vision-language representations to the physical constraints of a specific robot. Remarkably, this stage alone enables the model to complete a variety of tasks, such as shirt folding and bin picking on in-domain objects.
Stage 3: The task-specific adaptation aims to make the robot master complex tasks. These tasks include completing long-horizon tasks and generalizing to novel objects.
While the model has learned from diverse robot behaviors and progressively developed dexterous skills for complex tasks, it faces limitations in very long-horizon, contact-rich scenarios such as folding crumpled shirts or executing continuous bin-picking. Prior approaches often rely on high-level policy models; for example, $\pi_{0}$ use SayCan to update instructions every two seconds. In contrast, we propose leveraging the innate reasoning abilities of the VLA model to directly guide robot action. We train the model using demonstrations annotated with substep reasoning — for instance, breaking “fold the shirt” into “smooth wrinkles,” “align sleeves,” and “secure folds” — enabling it to learn disentangled action representations that map language sub-instructions to precise motor primitives.
We evaluate DexVLA across diverse embodiments, including single-arm, bimanual, dexterous hand, and mobile bimanual robots, demonstrating its effectiveness on a variety of tasks. DexVLA achieves high success rates on many tasks without task-specific adaptation. For example, it achieves a near full score (0.92) in folding flattened shirts. It can also learn dexterous skills on novel embodiments with fewer than 100 demonstrations, such as pouring drinks with a dexterous hand and packing on a bimanual UR5e. Furthermore, our approach can complete complex, long-horizon tasks like laundry folding using only direct instructions, eliminating the need for a high-level policy model to decompose the task. Importantly, our model is pre-trained on only 100 hours of demonstration data and runs at 60Hz on a single Nvidia A6000 GPU, enabling cost-efficient training and fast inference.
In summary, our contributions are three-fold:
  • We introduce Vision-Language Diffusion Expert (DexVLA), a novel architecture that scales the action expert to facilitate cross-embodiment learning.
  • We present an embodied curriculum learning strategy tailored for efficient and effective training of DexVLA. Furthermore, we enhance the VLA model’s learning by incorporating demonstrations annotated with sub-step reasoning.
  • We conduct comprehensive experiments across multiple embodiments, demonstrating DexVLA’s adaptability to challenging tasks without task-specific adaptation, its ability to learn dexterous skills on novel embodiments with limited data, and its capacity to complete complex, long-horizon tasks using only direct language prompting.
II Related Work
Vision-language-Action models for robot control. Recent research has focused on developing generalist robot policies trained on increasingly expansive robot learning datasets [12, 15, 13, 18, 17, 24, 36, 9, 14]. Vision-language-action models (VLA) [23, 5, 33, 50, 7, 62, 60, 16, 2] represent a promising approach for training such generalist policies. VLAs adapt vision-language models, pre-trained on vast internet-scale image and text data, for robotic control [55]. This approach offers several advantages: leveraging large vision-language model backbones, with billions of parameters, provides the necessary capacity for fitting extensive robot datasets. Furthermore, reusing weights pre-trained on internet-scale data enhances the ability of VLAs to interpret diverse language commands and generalize to novel objects and environments. However, current VLA models do not specifically focus on learning dexterous robotic skills by leveraging the parameters of the underlying VLM. While a few works, such as $\pi_{0}$ [5] and TinyVLA [50], introduce external action experts to facilitate action learning, their training pipelines still rely on the entire model. Another challenge is that even advanced methods like $\pi_{0}$, despite being capable of completing highly dexterous and long-horizon tasks, require the assistance of a high-level policy, such as SayCan [1], to decompose tasks into sub-goals. This allows the VLA to complete sub-tasks sequentially. We aim to integrate this high-level planning capability directly into the model itself by training each component of the network with data annotated at the sub-step level. Consequently, our method can complete complex tasks, like laundry folding, without requiring an external high-level policy, making the entire framework more end-to-end and demonstrating significant potential.
Diffusion models. Diffusion models [11, 32, 19] have emerged as the dominant approach in visual generation. The Diffusion Policy [8] successfully applies the diffusion model to robot learning, demonstrating its ability to model multimodal action distributions. Subsequent research has further developed the Diffusion Policy [61, 48, 34, 38, 41, 42, 3, 4, 10, 25, 37, 47, 29] by applying it to 3D environments [22, 57, 56, 54], scaling its capabilities [63], improving its efficiency [21, 34], and incorporating architectural innovations. There are a number of works investigating the usage of diffusion VLA [50, 5, 49]. Although existing models achieve strong performance and generalization on diverse tasks, they predominantly rely on the capabilities of pre-trained vision-language models. This work proposes a paradigm shift towards the diffusion module, demonstrating that a newly designed diffusion-based action expert, coupled with a novel training strategy, enables VLA models to learn from data more efficiently and effectively.
Figure 1: DexVLA architecture and embodied curriculum learning. Our model employs a three-stage training process. Stage 1 (left) trains the Diffusion Expert independently, without the VLM. Stages 2 and 3 (middle) integrate the Diffusion Expert with a VLM, discarding the visual and language components within the expert. The Diffusion Expert (right) uses multiple heads for cross-embodiment learning.
III Method
This section gives a detailed introduction to our model architecture and training strategy. The overview is presented in Figure 1.
III-A Model Architecture
Our DexVLA model is primarily based on a transformer language model backbone. Following the common framework of VLM models, we employ image encoders to project the robot’s image observations into the same embedding space as the language tokens. For multiple camera views, these visual tokens are concatenated. The VLM component generates two outputs: reasoning tokens and action tokens. The action tokens are passed through a projection module, consisting of two linear layers with LayerNorm. This module is analogous to the connectors designed in vision-language models like LLaVA [28], and serves to transform the VLM’s embedding space to align with the input requirements of the action expert. The reasoning tokens are injected into the policy model using FiLM layers, which scale and shift the parameters of the projection layers within the policy. Consequently, the model can autonomously generate reasoning and leverage this reasoning within the diffusion expert to guide action generation.
Following the standard late fusion VLM recipe [27, 28], image encoders embed the robot’s image observations into the same embedding space as language tokens. We further augment this backbone with robotics-specific inputs and outputs — namely, proprioceptive state and robot actions.
Building diffusion expert. Since action experts dominate the learning process of robot’s action, it is essential to design good neural architecture for better visuomotor policy learning. We utilized the Scale Diffusion Policy (ScaleDP [63]), a variant of the Diffusion Policy in Transformer architecture, where the largest version of ScaleDP is up to 1B parameters. However, the naive ScaleDP is not designed for cross-embodiement pre-training. We design a multi-head output to enable pre-training on ScaleDP with various robot configurations. Each head is responsible for a single robot configuration. This setup is similar to Octo.
Training objectives. Given a batch of input sequences, the overall training loss is defined as a weighted combination of the diffusion loss ($L_{diff}$) and the next-token prediction (NTP) loss ($L_{ntp}$):
$L=L_{diff}+\alpha L_{ntp},$ (1)
where $\alpha$ is a hyperparameter controlling the relative weight of the NTP loss term, $L_{ntp}$. Empirically, without the stage 1 training, the magnitude of $L_{ntp}$ is consistently an order of magnitude smaller than that of $L_{diff}$. Yet, after we adopted stage 1 training to warm up the policy learning ability of diffusion experts, we found that the magnitude of $L_{ntp}$ becomes slightly smaller than that of $L_{diff}$. Therefore, we set $\alpha=1$ in all experiments. This adjustment ensures that both the diffusion and next-token prediction objectives are comparably weighted during training, thus enabling the model to learn effectively from both action prediction and language understanding tasks. The following section provides a comprehensive introduction to the training strategy, detailing how it leverages the capabilities of our model architecture.
III-B Embodied Curriculum Learning
Curriculum learning is an training strategy where a system learns tasks in a progression from simple to complex, mirroring how humans acquire skills. Our three-stage training strategy implements an embodied curriculum, where the agent first learns generalizable motor skills from cross-embodiment data (Stage 1), then adapts to its specific physical form (Stage 2), and finally refines task-specific behaviors (Stage 3). This mirrors human skill acquisition, where foundational abilities (e.g., grasping) precede specialized expertise (e.g., folding clothes). By decoupling embodiment-agnostic and embodiment-specific learning, we reduce data needs by 60% compared to end-to-end training.
A well-designed training strategy is critical for optimizing deep neural networks. Approaches that align with a network’s inherent training dynamics ensure more efficient and effective data utilization. Our framework, DexVLA, targets general robotic control by integrating a Vision-Language Model (VLM) with a diffusion expert. Leveraging its modular architecture—which combines two distinct components—we propose a three-stage training strategy that systematically addresses: (1) learning dexterous manipulation skills to enable the model to complete complex tasks; and (2) cross-embodiment learning to adapt the model to diverse robotic platforms.
III-B1 Stage 1: Cross-Embodiment Pre-training
The vision-language-action models can be viewed as a composite of two distinct components. At the top of the architecture lies the vision-language model (VLM), which processes both visual input and language instructions, mapping them into a shared embedding space. This shared space is pre-trained using internet-scale data, enabling a wide range of capabilities, including language understanding, multimodal understanding, and various other vision-text tasks. However, despite its extensive training, the VLM lacks the ability to physically interact with diverse objects in real-world environments.
To effectively pre-train the action expert, we leverage all available data while temporarily decoupling it from the VLM component. This allows us to focus on developing a robust action generation capability independent of language grounding. While CNN architectures, such as ResNet, are commonly used as visual encoders, the visual encoders within VLM models typically employ a Transformer architecture. Therefore, for ease of alignment in subsequent stages, we utilize a ViT architecture for visual encoding. For language embedding extraction, we use DistilBERT [39]. The resulting language embeddings are then integrated into the model using FiLM layers, consistent with previous work. A key insight for pre-training the diffusion expert is the necessity of decomposing long-horizon tasks (e.g., bussing tables, laundry folding) into sub-tasks. These tasks, often spanning beyond 2 minutes, prove challenging for the diffusion expert to learn effectively from a single language instruction. Therefore, we annotate sub-step instruction within these long-horizon tasks to provide a more structured learning signal. Pre-training with sub-steps is crucial for strong performance. We demonstrate the importance of sub-step reasoning in the ablation study (see Appendix). Our empirical observations show that VLA trained without this pre-training frequently skip critical steps in very long tasks. Sub-step annotations are typically provided every five seconds of the demonstration.
Pre-training the large-scale robotic data on the diffusion-based action expert, rather than the entire VLA, offers several advantages. First, our diffusion expert has a significantly smaller parameter count compared to VLAs (1 billion vs. 3 billion), resulting in much faster training. Quantitatively, pre-training on the diffusion expert is five times faster than pre-training on the full VLA model.
Figure 2: Example of direct prompting for long-horizon tasks. The figure shows three tasks, laundry folding (top), dryer unloading (middle), sorting (bottom). Our DexVLA breaks down raw instructions into sub-steps automatically. Success in these tasks necessitates not only dexterity but also the capacity to decompose direct prompts into implicit multi-step reasoning and to comprehend the visual context.
III-B2 Stage 2: Embodied-Specific Alignment
While Stage 1 learns basic motor skills from cross-embodied data, this cross-embodiment learning can potentially compromise performance on the target embodiment, making it unsuitable for real-world deployment. Stage 2 addresses this by training the model with embodiment-specific data, aligning the abstract vision-language representations from the VLM with the diffusion expert. Therefore, we filter the dataset to include only embodiment-specific data, ensuring each sample involves a single embodiment. Mirroring techniques employed in vision-language models like LLaVA [28, 27], this stage focuses on aligning the target embodiment’s action space with its corresponding camera views and accompanying language instructions. Specifically, we jointly train the VLM model, the projection layer, and the diffusion expert on this embodiment-specific data, while freezing the VLM’s visual encoder. This joint training allows the diffusion expert to effectively ground the high-level vision-language understanding from the VLM in the specific motor control space of the target robot. Following Stage 2 training, we observe that the model exhibits proficiency in performing a range of tasks on the target embodiment, such as shirt folding, demonstrating the effectiveness of the embodiment-specific training.
Figure 3: Our experiment includes various robot types: bimanual UR5e, Franka, bimanual AgileX, and Franka with dexterous hands.
III-B3 Stage 3: Task-Specific Adaptation
This phase refines the model’s ability to execute downstream tasks skillfully and fluently, analogous to the post-training stage in large language models where the model is fine-tuned on domain-specific data. For simpler, less generalization-dependent tasks, such as shirt folding, table bussing, or bin picking with trained objects, task-specific training is unnecessary as the model already performs well. However, complex, dexterity-demanding tasks require the model to learn fine-grained, context-dependent actions. Therefore, effective post-training relies on a high-quality dataset of expert demonstrations exhibiting consistent and fluent task execution strategies focused on behaviors that promote successful task completion. It is worth noting that we utilize sub-step annotated language data in both Stage 2 and Stage 3. However, instead of directly using these sub-step reasoning as instructional input, we employ them as intermediate language output, compelling the model to learn and generate these sub-step language descriptions. This approach has proven highly effective, enabling our model to perform complex, long-horizon tasks such as laundry folding. While other VLAs, like $\pi_{0}$ [5], can also complete such tasks, they rely on high-level policy models like SayCan [1] to identify the task state and provide next-step instructions. In contrast, our framework leverages the VLM backbone as an implicit high-level policy model. This allows the model to internally interpret the task’s state and inject this understanding into the policy to guide action generation, eliminating the need for an external high-level policy module.
III-C Implementation Details
In our experiments, we employ the Qwen-2-VL 2B [46] model as our base vision-language model. We utilize up to three camera views. Our diffusion expert comprises 1 billion parameters. Two MLP layers, each followed by LayerNorm, connect the vision-language model to the action experts. We use FiLM to inject the final embedding from the vision-language model into the diffusion expert, applying it to each of the two MLPs. During Stage 1 training, we use our collected dataset. Stage 2 training is restricted to data sharing the same embodiment. Post-training is performed selectively, only on tasks where it is explicitly mentioned. Our pretrained dataset has a total of 100 hours of data with 91 different tasks. The detailed data allocation is referred to the Figure 4. For clarity, we report data quantity in hours rather than trajectories, as the duration of tasks varies significantly (i.e., laundry folding takes considerably longer than other tasks). This method of data quantification is consistent with that used by $\pi_{0}$. The detailed hyperparameters are listed in Table I. Specific details on implementation can be found in the Appendix.
TABLE I: Detailed hyperparameters for training Dex-VLA.
Hyperparameters Stage 1 Stage 2 Stage 3
Learning rate 1e-4 2e-5 2e-5
LR scheduler Constant Constant Cosine
Weight decay 0.0 0.0 0.0
Optimizer AdamW($\beta_{1}=0.9$,$\beta_{2}$=0.95)
Training epochs 5 5 5
Data Cross-embodied Embodied-specific Task-specific
Figure 4: Overview of our dataset for stage 1 training. Our dataset comprises approximately 100 hours of collected data spanning 91 distinct tasks. The majority of this data was collected using two robot platforms: the Agilex (ARX arm) (42.7%) and the single Franka Emika robot (34.7%). The “ARX arm” and “PIPER arm” represent two distinct robotic arm configurations, both featuring six degrees of freedom (6-DoF) but differing in their kinematic structures and operational characteristics.
III-D Robot System Setup
Our evaluation is conducted on four different robot configurations across 10 tasks. These setups are summarized in the following list and visualized in the accompanying figure:
  • Franka with gripper. A Franka Emika robot with 7 degrees of freedom, equipped with a Robotiq parallel jaw gripper. Data is collected at 15Hz.
  • Franka with dexterous hand. An Inspired Dexterous Hand mounted on a Franka Emika robot. The camera setup is identical to the gripper version, with a total of 12-dimensional configuration. Specifically, SE(3) end-effector pose (3D position + 3D orientation) plus 6-dimensional hand joint space. Data is collected at 15Hz.
  • Bimanual UR5e. Two UR5e robots, each with a Robotiq parallel jaw gripper and a wrist-mounted camera. A top-down camera is positioned between the two arms. This setup has a total of three camera views and a 14-dimensional configuration and action space. Data is collected at 10Hz.
  • Bimanual AgileX. Two 6-DoF AgileX arms, each with a wrist-mounted camera and a base camera. This setup has a 14-dimensional configuration and action space, supported by three cameras in total. Data is collected at 10Hz.
IV Real World Experiments
Figure 5: Examples of tasks without task-specific adaptation. We assessed our model’s performance after stage 2 training using three tasks: bin-picking easy (top), shirt folding (middle), and table bussing easy (bottom). Successful execution of these tasks necessitates a combination of dexterous manipulation skills.
Figure 6: Results on tasks without task-specific adaptation. We compared our model against Octo, OpenVLA, and Diffusion Policy. Performance was evaluated across 10 trials for each model, with scores averaged across these trials.
IV-A Evaluating Model without Task-Specific Adaptation
This section evaluates the model’s performance before task-specific adaptation (Stage 3). The evaluated tasks, visualized in Figure 5, all use the model with one set of parameters.
  • Shirt folding (Bimanual AgileX): The shirt is placed flattened on the table, and the robot is asked to fold a t-shirt. We evaluate two shirts, a yellow shirt of medium size and a blue shirt of large size.
  • Bin picking easy (Franka with gripper): The model needs to pick up all items from the right panel to the left tray. All items are seen in the dataset.
  • Bussing table easy (Bimanual AgileX): The robot must clean a table, place dishes and cutlery in a bin, and trash into a trash bin.
These tasks vary significantly in trajectory length and complexity, with some requiring high dexterity and intricate manipulation (e.g., shirt folding). We benchmark our approach against OpenVLA [23], a state-of-the-art 7B-parameter VLA model pre-trained on the Open X-Embodiment (OXE) [31] dataset, and Octo [30], a compact 93M parameter model that employs a diffusion-based policy for action generation. We used the open-sourced pre-trained weights for these two models. All baselines are fine-tuned on the same dataset and for the same number of epochs as our Stage 2 training, ensuring a fair comparison. We also compare to the Diffusion Policy [8], a strong baseline. Notably, neither Octo nor OpenVLA has previously demonstrated success on tasks of this complexity. While proprietary VLA models like $\pi_{0}$ and RT-2 [7] represent strong competitors, their closed-source implementations and reliance on private training data preclude direct empirical comparison.
Following $\pi_{0}$, we use a normalized score averaged over 10 episodes per task and method as our evaluation metric. Detailed scoring rubrics for each task are provided in the Appendix. As shown in Figure 6, DexVLA significantly outperforms all baselines on all tasks without task-specific adaptation. Notably, baseline methods, including OpenVLA, Octo, and Diffusion Policy, struggled to complete any steps of the shirt folding task, highlighting its complexity. In contrast, DexVLA achieves a 0.92 point on shirt folding without any task-specific adaptation. A similar phenomenon is observed in the bin picking and table bussing tasks. While these challenging tasks sometimes see limited success from the baselines, their overall scores remain low. DexVLA, however, achieves substantially better performance on these tasks.
Figure 7: Example of tasks for learning dexterous skills on new embodiment. We evaluate our model on two new embodiments with packing (top) and drink pouring (bottom) tasks, which are not included in stage 1 & 2 train data.
Figure 8: Results on learning dexterous skills from new embodiment. We evaluated our model with four three baselines: Diffusion Policy, Octo, and OpenVLA. Diffusion Policy is directly trained on these novel tasks from scratch.
IV-B Learning Dexterous Skills on New Embodiment
This section evaluates our model’s ability to learn dexterous skills on the new embodiments as shown in Figure 7. Specifically, the new embodiments are not involved in either Stage 1 or Stage 2 training. We aim to demonstrate the effectiveness of our proposed framework in acquiring new skills quickly on any new embodiment without the necessity of pre-training.
  • Drink pouring (Franka with dexterous hand): The drink is placed on the right of the table and a cup is placed on the left. The robot needs to grab the drink and pour it into the cup. This task includes 100 demonstrations.
  • Packing (Bimanual UR5): The robot is asked to pick up objects on both sides and place them into the box for packing. This task includes 100 demonstrations.
These tasks involve two novel robotic systems absent from training data: 1) A Franka arm integrated with a dexterous hand have 12 DoF, serves as a more complex robotic system than a simple gripper, posing a greater challenge for the model to learn. 2) A bimanual UR5e system featuring humanoid-inspired kinematic design and its articulation fundamentally differs from conventional dual-arm platforms like the AgileX bimanual robot. We evaluate our approach against three baseline methods: Diffusion Policy, Octo, and OpenVLA, the same baseline as in the previous section. This section aims to validate the adaptability of our pre-trained model to new embodiments and tasks. To this end, we directly fine-tune our Stage 2 pre-trained model on the novel tasks. For OpenVLA and Octo, we employ their publicly available checkpoints pre-trained on the OXE dataset. The Diffusion Policy — a method specialized for learning dexterous tasks from limited data — is trained exclusively on the two novel tasks. All methods are trained on individual tasks, and to ensure fairness, each baseline undergoes the same number of training epochs.
Figure 8 compares the performance of the methods on two novel tasks. For each method and task, we report the averaged scores over 10 trials (detailed scoring criteria are provided in the Appendix). DexVLA achieves an average of 0.90 point across two tasks, while OpenVLA and Octo struggle. DexVLA significantly outperforms Diffusion Policy, achieving a substantial performance lead. These results highlight DexVLA’s ability to efficiently adapt to new embodiments and master complex skills with only 100 demonstrations. These results are particularly meaningful as our method outperforms both extensively pre-trained VLA models (OpenVLA) and methods specifically designed for learning new tasks (Diffusion Policy).
Figure 9: Average scores on tasks requiring stage 3 training. We compared our model against two baselines: Octo and OpenVLA. Averaging scores over 10 trials, our method significantly outperformed both baselines across all tasks. Note that sorting was not included in the pre-training data.
IV-C Complex Long-Horizon Tasks with Direct Prompting
In this set of experiments, we tackle a range of challenging multi-stage tasks via a combination of task-specific post-training and self-generated reasoning capability. For some of these tasks, data is present in pre-training, but fine-tuning is required to attain mastery. For some, no data is present in pre-training. The tasks in this evaluation, shown in Figure 2, are:
  • Laundry folding (Bimanual AgileX): This task requires a static (non-mobile) bimanual system to fold articles of clothing. The clothing items start in a randomized crumpled state in a bin, and the goal is to take out the item, fold it, and place it on top of a stack of previously folded items. The randomized initial configuration of the crumpled laundry presents a major challenge since the policy needs to generalize to any configuration. This task is present in pre-training.
  • Dryer unloading (Bimanual AgileX): Here, the AgileX mobile robot has to take the laundry out of a dryer and place it into a hamper. This task is present in pre-training.
  • Sorting (Franka with gripper): The model needs to pick up all items from the right panel, and place them in the correct subsection of the left tray. The left tray is divided into four subsections. All new objects belong to the same category as these four sections. This task includes 200 demonstrations. This task is not in the pre-training data.
  • Bin picking hard (Franka with gripper): The model needs to pick up all items from the right panel to the left tray. Unlike the easy version, all objects are new and only present at test time. This task is present in pre-training.
  • Bussing table hard (Bimanual AgileX): The robot must clean a table, place dishes and cutlery in a bin, and trash into a trash bin. Unlike the easy version, all objects are new. In particular, we use dishes with unseen colors and trash with different appearances. This task is present in pre-training.
These tasks involve extended-horizon challenges. For instance, laundry folding requires more than 2 minutes to collect a single episode, and the soft, deformable fabric of clothing generates numerous unseen shapes and states, posing significant challenges for recognition and task completion. In the sorting task, the model must pick up 5–8 randomly placed objects in a cluttered scene and relocate them to predefined target positions.
We evaluate all models using averaged scores over 10 trials (detailed scoring criteria are provided in the Appendix). Conduct comparisons on these tasks is challenging due to the scarcity of prior models capable of operating at this scale. While $\pi_{0}$ represents the state-of-the-art in laundry folding, as mentioned before, its closed-source implementation and reliance on proprietary training data prevent direct benchmarking. Consequently, we compare our method only against OpenVLA and Octo. As demonstrated in Figure 9, DexVLA consistently surpasses all baselines and achieves a 0.8 point in the dryer unloading task. Even facing the most complicated task, laundry folding, our method attains a 0.4 point which proves its potential in handling extremely complicated scenarios. Through comprehensive evaluations, DexVLA demonstrates the capability to perform versatile, complex, and dexterous tasks via the combination of our proposed embodied curriculum learning method, scalable diffusion expert, and our novel VLA framework. We argue that DexVLA provides a promising approach for constructing VLA systems that effectively handle heterogeneous robotics data and master intricate manipulation tasks.
V Conclusion
This work proposes DexVLA, a novel architecture that leverages vision-language models to learn semantic information and employs a billion-parameter diffusion expert to learn robust and generalizable visuomotor policies. We introduce an embodied curriculum learning strategy, enabling the network to progressively learn from embodiment-agnostic motor skills to complex, embodiment-specific dexterous skills through three training stages. Furthermore, we incorporate sub-step reasoning, allowing the model to perform very long-horizon tasks without relying on a high-level policy model. Our method is evaluated from multiple perspectives, including its ability to perform complex tasks without task-specific adaptation, fine-tune on new embodiments with limited data, and execute extremely complex, long-horizon tasks without the assistance of a high-level policy model. Across all tasks, our method outperforms state-of-the-art approaches such as OpenVLA, Octo, and Diffusion Policy. We also conduct ablation studies to verify the effectiveness and efficacy of our proposed method.
Limitations and future work. Our method has several limitations and areas for improvement. First, we did not observe a significant transfer of learning from data collected on other embodiments to a specific target embodiment. Second, we observed that complex tasks, such as laundry folding, strongly depend on accurate action recovery and the ability to recognize the correct object state within the image. Prior work, $\pi_{0}$, addresses this by using SayCan, where a high-level policy model frequently assesses the object state and provides updated language instructions to the low-level vision-language-action model. Our current method, relying solely on the vision-language component, struggles with this aspect. A potential solution is to design a mechanism for explicit state checking and integrate it into our end-to-end model, which we plan to investigate in future work. Similar to observations reported in $\pi_{0}$, we also found that models trained with stage 3 exhibit reduced cross-task generalization compared to models trained solely through stages 1 and 2. However, complex tasks like laundry folding necessitate stage 3 training to effectively learn the required actions and object states for successful completion. Addressing this trade-off between task-specific learning and generalization is crucial for enabling models to perform complex tasks effectively without relying on post-training.
VI Acknowledgement
We would like to thank Yanjie Ze for the valuable discussion. We also like to thank Xiaoyu Liu for contributing to data construction, and the efforts of Xiaoyu Liu, Zhongyi Zhou, Chengmeng Li, Yefei Chen, Yaxuan Li, and Peiting Zhao in data collection.